# The Scale for the Expansion of the Universe: From Local Structures to Cosmology

### v1

@@ V Curvature Phase Space and Discussion
For spherically symmetric objects, binary motion, and dwarf galaxies orbiting main halos, dimensionless parameters quantify the ratio between the curvature arising from mass attraction and the curvature arising from the expansion of the universe. These parameters include the spherical density $\Delta_{c}$, which represents the ratio between the enclosed matter density and the cosmological density; the binary system parameter $\kappa$, defined as the ratio of the orbital period squared to the Hubble parameter squared; and the galaxy group and cluster parameter $\kappa$, which is the ratio of the turnaround radius to the distance determined by $\Lambda$. These dimensionless constants emerge naturally in the curvature scalars and characterize the interplay between local gravitational attraction and the expansion of the universe.
@@
The curvature scalars serve as key indicators of this transition, emerging naturally from fundamental theory and providing a unified framework for understanding the influence of dark energy. Fig. (3) illustrates the curvature phase space for the potential $\Phi$ versus the scalar $\sqrt{\mathcal{K}}$ in a de-Sitter universe. The Solar System and S-stars are positioned much higher in the phase space, reflecting strong local gravitational effects, while galaxy groups and clusters lie at the transition to local expansion, where DE dominates. This contrast highlights the scale-dependent influence of dark energy: smaller systems remain dominated by local gravity, whereas larger structures exhibit curvature scalars that approach their cosmological values, indicating the dominance of DE. By comparing systems of varying sizes, we have shown that DE dominates in larger structures, such as galaxy groups and clusters, where the curvature scalars align with cosmological predictions. This phase-space analysis provides a clear and unified framework for understanding the interplay between local gravitational dynamics and the large-scale expansion of the universe driven by dark energy.
The cases studied in this research are based on idealized cases. To probe the changes in galaxy groups and cluster with more complicated environment, we use N-body simulations. In that way, the upper panel of Fig (2) shows that the theoretical model fit to the simulation.

### v2

@@ V Discussion
From the McV spacetime, the dimensionless parameter $\kappa$ emerges, which marks a critical transition point across all systems studied. Of particular significance is the emergence of a universal dimensionless parameter, $\kappa$, from the curvature scalars of the McVittie spacetime. Despite originating from different physical contexts—be it the dynamics of galaxies, the collapse of overdensities, or the orbits of binaries—this parameter consistently delineates the frontier where DE’s repulsive influence begins to dominate over local gravitational attraction. For galaxy groups and clusters, $\kappa\equiv(r/r_{\Lambda})^{3}$ (from Eq. 7). The condition $\kappa=1$ defines the maximal turnaround radius $r_{\Lambda}$ (Eq. 24), the point of zero acceleration where a test particle is precisely balanced between the inward gravitational pull and the outward push of cosmic acceleration. This represents the absolute limit of the gravitational influence of a structure. Our analysis shows that the observed turnaround radii of local structures align closely with this $\kappa\sim 1$ transition, as evidenced by their position on the curvature phase space (Fig. 4) and their distribution relative to the zero-gravity surface (Fig. 3). This indicates that these massive structures exist at the very threshold of being unbound by DE.
In the spherical collapse model, the same parameter $\kappa$ represents the density contrast $\Delta_{c}=\rho_{m}/\rho_{\Lambda}$ (Eq. 7). Here, $\kappa=1$ ($\rho_{m}=\rho_{\Lambda}$) signifies the scale where local matter density equals the homogeneous DE density. While virialization itself occurs in the deep potential wells where $\kappa=\Delta_{c}\gg 1$, the $\kappa=1$ threshold defines the larger cosmological boundary within which the collapse process is embedded. The presence of DE modifies the critical overdensity for collapse from its Einstein-de Sitter value Pace et al. (2017), demonstrating that $\Lambda$ influences structure formation well before the $\kappa=1$ boundary is reached.
For binary systems, $\kappa$ takes the form $(T_{\text{Kep}}/T_{H})^{2}$. The condition $\kappa=1$ corresponds to an orbital period equal to the Hubble time ($T_{\text{Kep}}=T_{H}$), representing a fundamental upper limit on the orbital period of any gravitationally bound system in an accelerating universe. No known binary system approaches this limit; the longest-period systems, such as the Milky Way-Andromeda binary ($T_{\text{Kep}}\approx 0.62T_{H}$, $\kappa\approx 0.38$), remain in the gravity-dominated regime ($\kappa<1$) but already exhibit measurable effects from $\Lambda$ Benisty et al. (2024).
This multi-scale analysis reveals a consistent narrative: while the specific definition of $\kappa$ is context-dependent, the $\kappa\sim 1$ transition universally marks the scale where a system’s dynamics are no longer dictated solely by local gravity but are fundamentally reshaped by the cosmic acceleration driven by DE. Our formalism, based on the curvature invariants of the McVittie spacetime, provides a powerful and consistent geometric language to identify and characterize this transition across the vast range of scales in the cosmos.
@@
The curvature scalars serve as key indicators of this transition, emerging naturally from fundamental theory and providing a unified framework for understanding the influence of DE. Fig. (4) illustrates the curvature phase space for the potential $\Phi$ versus the scalar $\sqrt{\mathcal{K}}$ in a de Sitter universe. The Solar System and S-stars are positioned much higher in the phase space, reflecting strong local gravitational effects, while galaxy groups and clusters lie at the transition to local expansion, where DE dominates. This contrast highlights the scale-dependent influence of DE: smaller systems remain dominated by local gravity, whereas larger structures exhibit curvature scalars that approach their cosmological values, indicating the dominance of DE. By comparing systems of varying sizes, we have shown that DE dominates in larger structures, such as galaxy groups and clusters, where the curvature scalars align with cosmological predictions. This phase-space analysis provides a clear and unified framework for understanding the interplay between local gravitational dynamics and the large-scale expansion of the universe driven by DE.
The cases studied in this research are based on idealized configurations. To extend the analysis to more realistic environments, such as galaxy groups and clusters, we employ N-body simulations. As shown in Fig. 2, the theoretical predictions align well with both the simulation results and the observational data of the LG. The McV framework provides an exact solution of Einstein’s equations for a single point-like mass embedded in an expanding FLRW background. This construction relies on spherical symmetry and the absence of energy exchange between the local source and the cosmological fluid. These assumptions, however, restrict its applicability: once multiple localized masses are considered, isotropy and spherical symmetry are broken, and no straightforward generalization of McV exists. Consequently, McV should be regarded as an idealized model, valid only for isolated systems. To overcome these limitations, we turn to cosmological simulations such as IllustrisTNG, which capture the complexity of realistic cosmic environments. The agreement between the simulated Hubble flow around large-scale structures and the predictions of the McV picture confirms the robustness of our analytical framework despite its idealizations. This approach not only clarifies the conditions under which DE dominates local dynamics but also offers a geometric perspective for interpreting observations of cosmic expansion.